OpenStudy (selena2345):

In which direction does the parabola x+9y2=1 open?

OpenStudy (aravindg):

First write it in the standard form.

OpenStudy (selena2345):

but idk how to

OpenStudy (aravindg):

x+9y2=1 9y2=1-x y2=-(x-1)/9

OpenStudy (selena2345):

ok

OpenStudy (aravindg):

Now do u knw the answer?

OpenStudy (selena2345):

yes

OpenStudy (aravindg):

Cool :)